OpenStudy (anonymous):

solve by finding square roots 5x^2-45=0

OpenStudy (anonymous):

5x^2-45=0 5x^2=45 x^2=45/5 x^2=9 x=+-3 :)

OpenStudy (anonymous):

thank you again :D your awesome

OpenStudy (anonymous):

;)